Three probelms, two I think I got correct and one I know I didn't.
---------------------------------------------------------------------
Add. Express your answer in simplest form.

(8/s)+(8/(s^2)) I came up with an answer, but it was not in the choices offered, so I haven't got a clue!!! :o(
----------------------------------------------------------------------
Simplify. (a^7b^2)^7
I came up with a^14b^9, correct?
-----------------------------------------------------------------------
Solve by using the quadratic formula. 5x^2 + x = 3
I came up with -1+-sqrt61/10, correct??
Thanks for all the help that is offered!!!


Answer by edjones(8007)

(8/s)+(8/(s^2))
(8s+8)/s^2
8(s+1)/s^2
.
(a^7b^2)^7
=(a^49b^14)
.
5x^2+x=3
5x^2+x-3=0
(-1+-sqrt61)/10
You are correct on this one.
congrats.
.
Ed